# Robust Offline Imitation Learning from Diverse Auxiliary Data

### v1

@@ D.1 Implementation
We largely follow the architecture and hyperparameters from DWBC [37] for fair comparison. The policy network is a 3-layer MLP with 256 hidden units and tanh outputs. The discriminator is a 4-layer MLP with 128 hidden units, with sigmoid outputs clipped to [0.1, 0.9]. In the PU learning objective 1, we replace the non-differentiable max with the softplus function to make the loss function differentiable. The Q-function network is an MLP of 3 layers with 256 units. All networks use ReLU activations and the Adam optimizer.
@@
The discriminator learning rate is set to $1e^{-4}$ and a cosine annealing scheduler is added. The policy and Q-function learning rate is set to $3e^{-4}$, with a policy weight decay of 0.005. The balancing factors $\alpha$ and $\beta$ are set dynamically based on the loss ratios. Considering the batch-wise BC loss on expert data to be $\lambda_{1}$, the batch-wise weighted BC loss on auxiliary data to be $\lambda_{2}$, and batch wise Q-function loss to be $\lambda_{3}$, then $\alpha=0.0013*\lambda_{1}/\lambda_{2}$ and $\beta=0.053*\lambda_{1}/\lambda_{3}$. The discount factor $\gamma$ is 0.5. The frequency of actor model update, $t_{freq}$ is set to 3 for all the environments.
@@
D.2 Dataset
All datasets are from D4RL [7], an offline IL benchmark. Expert trajectories for locomotion tasks are from $<$environment$>$-expert-v2. Expert trajectories for manipulation tasks are from $<$task$>$-expert-v1. Sub-optimal transitions are from $<$environment$>$-random-v1 (locomotion) and $<$task$>$-cloned-v1 (manipulation). Table 7 details the number of trajectories and transitions in each dataset. The column $\textrm{Transitions}^{\dagger}$ refers to the full D4RL datasets for the corresponding random transitions with total number of trajectories given in brackets.
@@ D.2 Dataset
Table 8: Performance on the Hopper task with ground-truth rewards.
Setting Method
ROIDA ROIDA w/ GT rewards
5 /  0 $84.63$ p m 16.01 $94.63$ p m 20.76
5 /  3 $86.66$ p m 21.94 $98.53$ p m 11.70
5 /  5 $88.45$ p m 8.46 $104.46$ p m 5.42
Avg. 86.58 p m 16.42 99.21 p m 14.11
Appendix E Additional comparisons
We present additional comparisons in Table 9 with another recent unpublished algorithm, BCDP [28], which tailors an offline IL algorithm specifically for scenarios where the auxiliary offline dataset contains no expert data. While this design can enhance performance when the auxiliary dataset lacks expert demonstrations, it can lead to suboptimal performance as the proportion of expert data in the auxiliary dataset increases. The results indicate the ROIDA significantly outperforms BCDP on all the seven environments.
Method Environment
Hopper Walker2D Ant Pen Door Hammer Relocate
Appendix F RLiable evaluation
In this section, we present evaluations using the rliable [2] framework for our algorithm ROIDA  and DWBC, which is the closest contender. The rliable framework aims to reliably evaluate performance with a limited number of runs by employing a rigorous evaluation methodology that accounts for uncertainty in results. It presents more robust and efficient aggregate metrics, such as interquartile mean (IQM) scores, to achieve small uncertainties in the evaluation outcomes.
We group the environments into two benchmarks, locomotion and adroit. Locomotion contains the Hopper, Halfcheetah, Walker2D and Ant environments and consists of 12 tasks (4 environments $\times$ [5/0, 5/3, 5/5]). Adroit contains the Pen, Door, Hammer and Relocate environments and consists of 12 tasks (4 environments $\times$ [50/0, 50/30, 50/50]). We divide the scores by 100 to obtain values in the [0, 1] range, as used in the rliable paper. Using this setup, we report the following performance metrics with 95% confidence intervals: 1) Median performance (higher better), 2) Mean performance (higher better), 3) IQM (higher better), and 4) Optimality gap (lower better).
Appendix G Limitations
In our particular framework, the reward estimation can be improved by an accurate choice of the hyperparameter $\eta$ by performing mixture proportion estimation [39, 23]. However, this is beyond the scope of our work. In order to avoid any assumption about the auxiliary data in our work, we have chosen $\eta=0.5$ which is an unbiased estimate. We also provide additional results with $\eta=0.3$ and $\eta=0.7$ in Table 10. Here, we obtain better results when $\eta$ is closer ($\eta=0.3$ is closer than $\eta=0.7$) to the true ratio between expert and suboptimal demonstration in the auxiliary dataset (0.01 for setting 5/0, 0.12 for setting 5/3 and 0.19 for setting 5/5). Since this true ratio is unknown, estimating it would be a problem in its own right, which could then be combined with our method.
Table 10: Performance on the Hopper task with varying $\eta$.
Setting Method
5 /  0 $88.12$ p m 14.93 $84.63$ p m 16.01 $82.40$ p m 20.76
5 /  3 $90.42$ p m 18.84 $86.66$ p m 21.94 $84.85$ p m 7.13
5 /  5 $91.02$ p m 7.97 $88.45$ p m 8.46 $86.45$ p m 18.61
Avg. 89.85 p m 14.62 86.58 p m 16.42 84.57 p m 14.39
Appendix H Broader impact

### v3

@@ A.2.1 Implementation
We largely follow the architecture and hyperparameters from DWBC (Xu et al., 2022) for fair comparison. The policy network is a 3-layer MLP with 256 hidden units and tanh outputs. The discriminator is a 4-layer MLP with 128 hidden units, with sigmoid outputs clipped to [0.1, 0.9]. In the PU learning objective 1, we replace the non-differentiable max with the softplus function to make the loss function differentiable. The Q-function network is an MLP of 3 layers with 256 units. All networks use ReLU activations and the Adam optimizer.
@@
The discriminator learning rate is set to $1e^{-4}$ and a cosine annealing scheduler is added. The policy and Q-function learning rate is set to $3e^{-4}$, with a policy weight decay of 0.005. The balancing factors $\alpha$ and $\beta$ are set dynamically based on the loss ratios. Considering the batch-wise BC loss on expert data to be $\lambda_{1}$, the batch-wise weighted BC loss on auxiliary data to be $\lambda_{2}$, and batch wise Q-function loss to be $\lambda_{3}$, then $\alpha=0.01*\frac{\lambda_{1}}{\lambda_{2}}*\frac{1}{7.5}$ and $\beta=0.01*\frac{\lambda_{1}}{\lambda_{3}}*\frac{1}{7.5}$. The additional factor of $\frac{1}{7.5}$ emphasizes the BC loss on expert data, and is adopted from previous work. The discount factor $\gamma$ is 0.5. The frequency of actor model update, $t_{freq}$ is set to 3 for all the environments.
@@
A.2.2 Dataset
All datasets are from D4RL (Fu et al., 2020), an offline IL benchmark. Expert trajectories for locomotion tasks are from $<$environment$>$-expert-v2. Expert trajectories for manipulation tasks are from $<$task$>$-expert-v1. Sub-optimal transitions are from $<$environment$>$-random-v1 (locomotion) and $<$task$>$-cloned-v1 (manipulation). Table 8 details the number of trajectories and transitions in each dataset. The column $\textrm{Transitions}^{\dagger}$ refers to the full D4RL datasets for the corresponding random transitions with total number of trajectories given in brackets.
@@ A.2.2 Dataset
A.3 RLiable evaluation
In this section, we present evaluations using the rliable (Agarwal et al., 2021) framework for our algorithm ROIDA  and DWBC, which is the closest contender. The rliable framework aims to reliably evaluate performance with a limited number of runs by employing a rigorous evaluation methodology that accounts for uncertainty in results. It presents more robust and efficient aggregate metrics, such as interquartile mean (IQM) scores, to achieve small uncertainties in the evaluation outcomes.
We group the environments into two benchmarks, locomotion and adroit. Locomotion contains the Hopper, Walker2D and Ant environments and consists of 9 tasks (3 environments $\times$ [5/0, 5/3, 5/5]). Adroit contains the Pen, Door, Hammer and Relocate environments and consists of 12 tasks (4 environments $\times$ [50/0, 50/30, 50/50]). We divide the scores by 100 to obtain values in the [0, 1] range, as used in the rliable paper. Using this setup, we report the following performance metrics with 95% confidence intervals: 1) Median performance (higher better), 2) Mean performance (higher better), 3) IQM (higher better), and 4) Optimality gap (lower better).
Figure 4: rliable evaluation on the Locomotion benchmark.
Figure 5: rliable evaluation on the Adroit benchmark.
A.4 Analysis on hyperparameters
In the main paper, we study the impact of $\eta$ and $\tau$. In this section, we analyze the effect of varying the weighting factors $\alpha$ and $\beta$, which control the contributions of the weighted BC loss and the Q-learning loss, respectively. The results are presented in Figure 6. We vary both hyperparameters across a range from 0.001 to 1.0. Based on the ablation study for the Hopper-5/5 scenario, we select 0.01 for both $\alpha$ and $\beta$ across all setups. These values are then dynamically scaled according to the loss balancing scheme outlined in Section A.2.
(a) Varying $\alpha\ (\beta=0.01)$
(b) Varying $\beta\ (\alpha=0.01)$
Figure 6: Performance on Hopper-5/5 as $\alpha$ and $\beta$ are varied.
Figure 7: Reward estimation accuracy on Hopper 5/5. By combining PU learning-based reward estimation with the filtering strategy, ROIDA effectively identifies the high-reward transitions.
A.5 Reward visualization
To understand the accuracy of the proposed reward learning and filtering strategy, we visualize the overlap between the distribution of high-reward transitions as estimated by ROIDA and the overall set of transitions for the Hopper environment. As shown in Figure 7, ROIDA correctly identifies the ground-truth high-reward transitions as expert transitions.